OpenStudy (jojo12g):

Subtract. –11 – (–15) A. –26 B. –4 C. 4 D. 26

OpenStudy (pinklion23):

what do you think

OpenStudy (alexandervonhumboldt2):

-(-)=+ so -11-(-15)=-11+15=15-11=???

OpenStudy (pinklion23):

C

OpenStudy (jojo12g):

i think b or a

OpenStudy (m_m13):

it's C. 4

OpenStudy (jojo12g):

thanks guys

OpenStudy (just_one_last_goodbye):

NEgative + negative = positive meaning 15 - 11 = 4

OpenStudy (ddos_dragon):

The answer is C.

OpenStudy (alexandervonhumboldt2):

@jojo12g -(-)=+ so -11-(-15)=-11+15=15-11=??? what is 15-11?

OpenStudy (ddos_dragon):

Make sure and medal whoever helped you out the best and close the question. Have a nice day.

OpenStudy (just_one_last_goodbye):

nice job @AlexandervonHumboldt2

OpenStudy (jojo12g):

i wish i can give medals to all of u guys